Given the functions f(x)=x^2 and g(x)=sqrt(x), evaluate the following:
(f+g)(7)=
(f+g)(18)=
(f+g)(-18)=

Answer by fractalier(6550)

Did this one earlier...
(f+g)(7)= f(7) + g(7) = 49 + sqrt(7)
etc...